OpenStudy (anonymous):

What is the first step in solving the equation 5x − 3 = 22?

OpenStudy (anonymous):

combining like terms

OpenStudy (anonymous):

5x - 3 = 22 5x = 22 + 3 5x = 25 5x = 25/5 = 5 x = 5

OpenStudy (anonymous):

remember to always try to combine like terms with equations that have variables

OpenStudy (anonymous):

x = 5 again lol

OpenStudy (anonymous):

but what you did 1st Add 3 to both sides of the equation. Divide both sides of the equation by −5. Divide both sides of the equation by −3. Subtract 3 from both sides of the equation ???????

OpenStudy (anonymous):

Add three to both sides first

OpenStudy (anonymous):

What I did first? Bring all numbers with coefficient x to one side, and all numbers to the other.

OpenStudy (anonymous):

then you do everything else because your main objective is to isolate x

OpenStudy (anonymous):

Brave ilikequadraticsokay! :D

OpenStudy (anonymous):

thanks :)

OpenStudy (anonymous):

brave? and your welcome(:

OpenStudy (anonymous):

Meant bravo* :P

OpenStudy (anonymous):

oohhh okay lol